### v1

# Scaling Laws Revisited: Modeling the Role of Data Quality in Language Model Pretraining

Abstract
Scaling laws for language model training traditionally characterize how performance scales with model size and dataset volume. Prior work has explored architecture variants and data treatments such as dataset filtering and noise injection in language model pretraining; however, these studies have not formalized data quality within a principled scaling law. We introduce a dimensionless data-quality parameter $Q$, and propose a quality-aware scaling law extending the Chinchilla framework to predict loss as a joint function of model size, data volume, and data quality. The law is motivated by an effective-sample-size and information-theoretic view of noisy or redundant corpora, and it admits two practical estimators for $Q$: (i) a corruption rate proxy and (ii) a deficiency measure. Through synthetic experiments in neural machine translation and autoregressive modeling—where we systematically control data quality via multiple levels of noise injection and coverage variation—we show that loss scales predictably with data quality and that higher-quality data can substantially reduce model size and hence compute requirements. Our results demonstrate a sublinear decay of effective data with quality and robustness to moderate data corruption; out-of-sample evaluations further validate the predictive form of the law. Unlike prior empirical analyses, our work establishes an explicit, generalizable law for data quality, offering concrete guidance for balancing data curation effort and model scale in large-scale pretraining.
1 Introduction
It is well understood in large-language model (LLM) training that both the amount of training data and its quality influence pretraining loss. Yet most existing scaling laws are formulated solely in terms of dataset size, remaining agnostic to quality [19, 16, 23, 35, 1, 8, 36, 41]. This leaves a gap between common intuition–“cleaner data trains better models”–and the quantitative laws that guide large-scale training.
The importance of such a quantitative framework is especially clear in specialized domains–business, scientific, or medical applications–where training corpora are limited in quantity and heterogeneous in quality. With the same compute budget, outcomes may differ drastically depending on how clean and representative the data are. There have been quite a number of studies to classify data quality among multiple “dimensions” [39, 14, 33, 22]. While the exact taxonomy varies, the core dimensions typically include: accuracy, completeness, consistency, timeliness, uniqueness, validity. The focus of these studies is usually datasets containing structured data that is clean and curated for a certain task or purpose. In contrast, when collecting data to pretrain large language models, large amounts of “found” data are collected and cleaned, usually by removing duplicate data and obviously “poor” quality data.
To reason about how such imperfect corpora affect performance, we need a simpler abstraction of quality. In this work, we introduce a single dimensionless parameter $Q\in(0,1]$ characterizing the usable information in a corpus. A value of $Q=1$ represents fully clean and representative data, while smaller $Q$ values reflect increasing corruption or redundancy. As discussed below (Section 3), $Q$ can be estimated via corruption rates or a more general deficiency measure, and serves as a proxy for the fraction of effective samples. We incorporate this abstraction into scaling laws through a term of the form $B/(D^{\beta}Q^{\gamma})$ (Equation 2), linking model performance directly to both the quantity and the quality of data.
Our main contributions are:
  • We propose a theoretically grounded quality-aware scaling law that augments the classical Chinchilla form by introducing a dimensionless quality parameter $Q$. The law predicts loss as $L(N,D,Q)=A/N^{\alpha}+B/(D^{\beta}Q^{\gamma})+E$, capturing the interplay between model size, data volume, and data quality.
  • We derive the scaling law from an effective sample size perspective and information‑theoretic arguments, and we provide two simple estimators for $Q$ based on corruption rate and data deficiency. We show that under natural assumptions these estimators lead to the form $D_{\mathrm{eff}}=D\,g(Q)$ with $g(Q)\approx Q^{\gamma}$, recovering the proposed law.
  • We conduct controlled experiments on neural machine translation and causal language modeling to illustrate the law’s utility. These results demonstrate that loss scales predictably with our quality parameter and that higher‑quality data can compensate for smaller models. This is particularly relevant for specialized models in many real-world domains, such as business, scientific or medical applications.
2 Related Works
Data quality in information systems.
Data quality has long been studied in information systems, where datasets are evaluated along multiple dimensions—accuracy, completeness, consistency, timeliness, uniqueness and validity—to ensure that structured records support downstream tasks. A consistent finding from this literature is that diverse imperfections degrade the utility of data and therefore motivate extensive curation practices [6]. Large-scale pretraining, however, relies on corpora assembled from the web. Here, quality varies widely across sources, and subtle forms of corruption or deficiency persist despite deduplication or heuristic filtering. Nonetheless, the key insight from information systems continues to hold: the quality of training data directly influences the performance of learned models.
Data quality in large-scale pretraining.
In the era of LLMs, quality has become a first-class design concern. Major corpora such as The Pile [15], RefinedWeb [32], and Dolma [34] highlight different philosophies of data construction, from assembling diverse sources to aggressively filtering and refining web text. Deduplication strategies [27] and contamination checks [30] have been shown to substantially affect evaluation reliability. Empirical analyses further suggest that careful filtering yields performance gains comparable to scaling compute, therefore implying the nontrivial role of quality in driving LLM performance.
Classical scaling laws and their limitations.
Alongside these developments, foundational studies on scaling laws formalized how performance scales with model size, dataset size, and compute. Kaplan et al. [25] showed that loss decreases predictably as models and datasets grow, while Hoffmann et al. [20] refined these results by introducing compute-optimal training. Despite their influence, these existing scaling laws generally assume that the underlying data is of fixed quality. In real-world corpora, noise, redundancy, or domain imbalance systematically distort scaling behavior, highlighting the need to explicitly account for quality.
Data quality and scaling laws.
Several recent works have begun to address this gap. Bansal et al. [5] study scaling laws for neural machine translation, and show that while test loss scales predictably with dataset size, architectural choices and moderate noise primarily shift the scaling curves without changing the exponent. Goyal et al. [17] argue that data filtering is not compute-agnostic: the value of higher-quality data depends on the available compute budget, since reusing small clean datasets leads to diminishing returns. Li et al. [28] assess data quality by comparing perplexity gaps between models of different sizes, theoretically linking their metric to inverse scaling laws. These studies highlight the importance of data quality but stop short of providing a unified predictive law that couples quality with model size and dataset size. By introducing a dimensionless quality parameter $Q$ into scaling laws, our work formalizes the influence of data quality and offers predictive guidance for building smaller, high accuracy models in specialized domains.
Beyond these empirical analyses, several recent studies refine scaling laws from complementary perspectives. Chang et al. [9] propose parameter-constrained scaling laws that combine text diversity and syntheticity, revealing that effective training tokens depend on both data quantity and diversity. On the theoretical front, Bahri et al. [3] present a unifying theory of neural scaling laws, distinguishing variance-limited and resolution-limited regimes and connecting scaling exponents to generalization error. Although these contributions deepen our understanding of scaling, they do not incorporate an explicit scalar quality parameter. Our work complements this literature by modeling data quality via $Q$, linking it to effective sample size and generalization theory, and providing a unified scaling law that accounts for quality.
3 Data Quality Measures
To incorporate data quality into scaling laws, we first need a formal definition of $Q$. In this section we introduce two approaches for mapping a dataset $\omega$ to a scalar $Q(\omega)\in(0,1]$, with larger $Q$ indicating higher quality. These definitions capture different sources of imperfection—token corruption versus more general deficiency—and will later allow us to connect $Q$ to effective sample size and generalization theory. The precise choice of $Q$ is application-dependent; the key requirement is that $Q$ degrades smoothly under corruption, providing a proxy for the usable information in a dataset.
3.1 Data Corruption Rate
One of the simplest measures of data quality is what is sometimes called the data corruption rate, which we will denote CR. We assume $0\leq\text{CR}<1$.
Definition 1.
Let $\omega$ be a dataset with data corruption rate CR. We define the data corruption rate data quality by
$Q(\omega)=1-\text{CR}.$
As an example, if we have a dataset consisting of $D$ tokens and $10\%$ are corrupted, then we would say that the $\text{CR}=10\%$ and the data quality $Q$ is $90\%$. To simplify our formulas, we assume that the data corruption rate is strictly less than $1$, so that $Q$ is strictly greater than 0. Note that standard sampling methods (e.g. [12]) can be used to estimate data corruption.
3.2 Data Deficiency Measures
We introduce a dimensionless quantity $\Delta$ call the data deficiency, which quantifies the lack of quality in a dataset $\omega$. We assume that the data deficiency has the following properties:
  1. 1.
    Positivity. $\Delta(\omega)\geq 0$, for all datasets $\omega$
  2. 2.
    Continuity. $\Delta(\omega)$ is a continuous function of the dataset $\omega$
  3. 3.
    Additivity. If $\omega_{1}$ and $\omega_{2}$ are independent datasets, and $\omega$ is the union of $\omega_{1}$ and $\omega_{2}$ than $\Delta(\omega)=\Delta(\omega_{1})+\Delta(\omega_{2})$.
Additive noise model. Data deficiency can be measured and quantified in many ways. One of the simplest models is to model data deficiency using an additive noise model, where each data measurement $\xi_{i}$, for $i=1$, $\ldots$, has the form
$\xi_{i}=x_{i}+\epsilon_{i},$
where $\epsilon_{i}$ is a noise term. We assume that $\epsilon_{i}\geq 0$, and that the data quality as measured by the data deficiency worsens as $\epsilon_{i}$ increases. Note that properties (1)—(3) above hold when we model data quality using the additive noise model. We now add a fourth assumption.
  1. 4.
    Maximum quality. In the additive noise model, we assume that the noise term $\epsilon_{i}$ captures all the noise, so that data quality is maximized when $\epsilon_{i}=0$, for all $i$.
Note that the data deficiency $\Delta$ can grow arbitrarily large with this model, so that $0\leq\Delta(\omega)<\infty$, for a dataset $\omega$.
Definition 2.
We define the data deficiency data quality $Q(\omega)$ of a dataset $\omega$ by
$Q(\omega)=\exp\left(-\Delta(\omega)\right).$
4 Quality-Aware Scaling Laws
A widely used empirical scaling law for LLMs (also known as the Chinchilla Scaling Law by Hoffmann et al. [20]) is
$L(N,D)=\frac{A}{N^{\alpha}}+\frac{B}{D^{\beta}}+E,$
where $N$ is the number of training parameters, $D$ is the number of training tokens, and $L(N,D)$ is the pretraining loss. This equation was estimated empirically from experimental studies involving over 400 LLM, with the number of parameters ranging from 70M to 16B and the number of training tokens ranging from 5B to 400B. Here $A$, $B$, $E$, $\alpha$, and $\beta$ are empirically estimated constants, with $E$ the minimal loss.
For example, one formulation studied in Hoffmann et al. [20] is:
$L(N,D)=\frac{406.4}{N^{-0.34}}+\frac{410.7}{D^{-0.28}}+1.69.$ (1)
In large-scale training, improvements such as filtering out low-quality data or introducing better training objectives have yielded improved performance beyond what naive scaling might predict. For instance, filtering training data to remove noise has been observed to effectively improve the scaling behavior, enabling models to achieve lower loss with the same compute. This indicates that not all tokens are equal: a billion high-quality tokens may be far more valuable than a billion noisy or redundant ones. The Chinchilla analysis already hinted that many models are undertrained given their size, pointing toward the need for either more data or better data. This motivates us to extend scaling laws with explicit parameters for data quality.
4.1 Effective Sample Size
We now formalize when data quality $Q$ leads to a scaling law of the form $D^{-\beta}Q^{-\gamma}$. The key idea is that poor data quality (e.g. data corruption or large data deficiency) reduces the effective sample size of the dataset. We formalize this with:
Definition 3.
Let $g$ be a monotone function $g:\![0,1]\to\mathbb{R}_{+}$ with $g(1)=1$. Then the effective sample size of $D$ associated with the link function $g$ is
$D_{\mathrm{eff}}\;:=\;D\,g(Q).$
Assumption 1 (Effective sample size factorization).
There exists a monotone function $g:[0,1]\to\mathbb{R}_{+}$ with $g(1)=1$ such that the expected excess loss satisfies
$L_{N}(D,Q)\;\approx\;\frac{B}{D_{\mathrm{eff}}^{\beta}}=\frac{B}{(D\cdot g(Q))^{\beta}}.$
where $L_{N}(D,Q)$ denotes the loss for a given parameter size $N$.
This assumption is consistent with classical results: in PAC learning with random classification noise of rate $\eta<1/2$, the sample complexity inflates by a factor $(1-2\eta)^{-2}$ [2, 26]; in regression with additive Gaussian noise of variance $\sigma^{2}$, effective samples scale with the signal-to-noise ratio $\mathrm{SNR}=1/\sigma^{2}$ [38]; and in information theory, channel capacity arguments show that mutual information is reduced by a multiplicative factor depending on the corruption level [10]. In the following two lemmas (proof deferred to the appendix), we show that the usable sample size can be written as $D_{\mathrm{eff}}=D\cdot g(Q)$ with $g(Q)$ well-approximated by a power law $Q^{\gamma}$ over practical ranges of $Q$:
  • $\gamma\approx 1$ for signal-to-noise ratio like token noise (Lemma 1)
  • $\gamma\approx 2$ for symmetric label noise (Lemma 2)
Lemma 1 (Effective sample size under additive Gaussian noise).
Consider i.i.d. observations of the form $y_{i}=f(x_{i})+\epsilon_{i},~\epsilon_{i}\sim\mathcal{N}(0,\sigma^{2})$, for $i=1,\dots,D$. Then the Fisher information contributed by each observation is proportional to $1/\sigma^{2}$, and the total Fisher information satisfies
$\mathcal{I}_{D}\;\propto\;\frac{D}{\sigma^{2}}.$
Equivalently, the $D$ noisy samples carry the same information as $D_{\mathrm{eff}}=D\cdot(1/\sigma^{2})$ noise-free samples.
We define the deficiency as
$\Delta=\ln\frac{\sigma^{2}}{\sigma_{0}^{2}},\qquad Q=e^{-\Delta}=\frac{\sigma_{0}^{2}}{\sigma^{2}}.$
where $\sigma_{0}^{2}$ is baseline noise (i.e. e.g. noise not modeled by $\Delta$). Without loss of generality, we may assume $\sigma_{0}:=1$. In this formulation, datasets noisier than the reference ($\sigma^{2}>\sigma_{0}^{2}$) have $Q<1$, while “clean” datasets ($\sigma^{2}=\sigma_{0}^{2}$) have $Q=1$. The effective sample size then scales as
$D_{\mathrm{eff}}=D\cdot Q^{\gamma}$
and taking $\gamma=1$ recovers the classical $D\cdot\mathrm{SNR}$ scaling.
Lemma 2 (Effective sample size under random classification noise).
Consider binary classification with labels $Y\in\{0,1\}$ and i.i.d. examples $(X,Y)$. Suppose labels are corrupted by random classification noise (RCN) with flip rate $\eta\in[0,1/2)$, i.e., we observe $\tilde{Y}=Y$ w.p. $1-\eta$ and $\tilde{Y}=1-Y$ w.p. $\eta$, independently of $X$. For any classification-calibrated surrogate loss admitting an unbiased noise-corrected estimator (e.g., logistic/hinge via the standard correction), the variance of the per-example corrected loss inflates by a factor proportional to $(1-2\eta)^{-2}$. Consequently, generalization bounds and excess-risk rates scale as if the number of clean samples were
$D_{\mathrm{eff}}=D\cdot(1-2\eta)^{2},$
i.e., the effective sample size is reduced by $(1-2\eta)^{2}$ under RCN.
Lemma 2 formalizes the intuition that symmetric label noise reduces usable information multiplicatively, yielding $D_{\mathrm{eff}}=D\cdot g(Q)$ with $g(Q)=(2Q-1)^{2}$. In other words, if we define data quality via the corruption rate as $Q=1-\eta$ (cf. Section 3), then $(1-2\eta)^{2}=(2Q-1)^{2}$. Over practical high-quality regimes $Q\in(1/2,1]$, the factor $(2Q-1)^{2}$ is well-approximated by a local power law in $Q$ with exponent $\gamma\approx 2$, so that an effective-sample model $D_{\mathrm{eff}}=D\cdot Q^{\gamma}$ with $\gamma\approx 2$ is justified for RCN.
4.2 Quality-Aware Scaling Law
We incorporate the data quality $Q$ and introduce the following quality-aware scaling law:
Definition 4 (Quality-Aware Scaling Law).
$L(N,D,Q)=\frac{A}{N^{\alpha}}+\frac{B}{D^{\beta}Q^{\gamma}}+E,$ (2)
where $Q$ denotes data quality and $\gamma$ is a empirically measured parameter.
Note that for the highest quality data with $Q=1$, this reduces to the standard Chinchilla scaling law, and as the quality of the data decreases, the loss increases, as is expected. On the other hand, as the quality of the data goes up, the amount of the data required to achieve the same loss goes down.
The following corollary formalizes how the proposed form arises under mild regularity conditions:
Corollary 1.
Suppose that $g(Q)$ is regularly varying at $Q=1$, i.e. $g(Q)=c\,Q^{\gamma}(1+o(1))$ as $Q\to 1$. Then the quality-aware scaling law (Definition 4) holds under Assumption 1.
We next give an information-theoretic perspective that leads to the same law:
Proposition 1 (Information-theoretic justification).
Let $X$ denote clean tokens, $\tilde{X}$ their corrupted versions produced by a memoryless channel $\mathcal{C}_{Q}$ parameterized by quality $Q\in(0,1]$, and let $Z$ be the learned representation used for prediction. Suppose corruption reduces usable information multiplicatively,
$I(\tilde{X};Z)\;=\;\rho(Q)\,I(X;Z),$
with $\rho(1)=1$, $\rho(Q)$ monotone, and locally $\rho(Q)\approx c\,Q^{\gamma}$ as $Q\to 1$. Here $I(X;Z)$ denotes mutual information as usual. If the data-dependent loss scales as [^1] This assumption is motivated by information-theoretic generalization bounds [42, 7], which relate generalization error to the mutual information between training data and learned representations or weights, as well as by the Information Bottleneck principle [37].
$L_{D}\;\propto\;\frac{1}{\big(D\cdot I(\tilde{X};Z)\big)^{\beta}},$
then the quality-aware scaling law takes the form
$L(N,D,Q)\;\approx\;\frac{A}{N^{\alpha}}+\frac{B}{D^{\beta}Q^{\gamma}}+E,$
after reparameterization of constants.
The proof of Proposition 1 is deferred to the appendix. For a binary symmetric channel with flip rate $\eta$, $\rho(Q)$ is proportional to the capacity $1-H_{2}(\eta)$ with $Q=1-\eta$; over practical ranges this is well-approximated by $Q^{\gamma}$ with $\gamma\approx 2$. In regression with additive Gaussian noise, $\rho(Q)$ is proportional to the signal-to-noise ratio, yielding $\gamma\approx 1$. Both cases recover the effective sample size view $D_{\mathrm{eff}}=D\cdot\rho(Q)$.
5 Experimental Studies
To establish the quality aware scaling laws, we train decoder only language models across two tasks, neural machine translation (NMT) and causal language modeling (CLM). To capture the effects of quality and dataset size we train on 3 different dataset volumes and 7 degrees of quality. We use subsets of Paracrawl v8 ([4]) and C4 ([13]) as our training data for the translation and language tasks respectively. See Table 1. We optimize cross-entropy loss averaged over the relevant context and report the same on held out test datasets to measure the effectiveness of our quality aware scaling law for predicting the loss on out of sample data (Table 2. We use the parametric loss fitting approach described in [20] to find our parameters using two methods, Least Squares and Huber [21]. The full details of the experiments, and model training details can be found in the appendix.
5.1 Task Description
Neural Machine Translation
Our first task is English to German translation using the Paracrawl v8 English to German dataset [4]. We filter the raw data with min hash deduplication, language ID filtering with a 0.8 threshold and length filtering ($\leq$ 258) leaving $\sim$ 101M sentence pairs. We use a 8L GPT Neo model with a hidden size of 1024 with approximately $\sim$ 133M params. We also train our own BPE tokenizer with a vocabulary size of 32000. We train 3 replicates for each combination of the 3 datasets of size 500K, 1M and 2M sentence pairs and 7 quality levels for a total of 63 experiment runs. We train using AdamW with a maximum learning rate of 5e-4 and cosine decay with a 20% warmup ratio. The full model configuration and training hyper-parameters are available in the appendix.
Causal Language Modeling
For the CLM objective we train our models on a subset of C4 (en) [13]. We run our pretraining experiments on a 8L Llama 3 [18] model with a hidden size of 512 and a context length of 2048. We do not use ROPE scaling. We employ random data truncation on the base dataset to sample within the context length. We also employ the pre-trained Llama-3.2-1B tokenizer. We sample datasets of three sizes; 100M , 1B and 10B tokens and train for a single epoch using fused AdamW [29] with a maximum learning rate of 1e-3 with cosine decay and a 10% warm-up ratio. We duplicate our experiment setting to match a total of 63 runs. The full model configuration and training hyperparameters are available in the appendix.
5.2 Synthetic Noise and Data Sampling Strategy
To simulate the different levels of data quality we add synthetic iid noise to the base dataset. The base dataset after pre-processing is assumed to have quality, $Q=1.0$. We vary the quality of a dataset by increasing/decreasing the fraction of samples that are perturbed with synthetic noise. For example, if $25\%$ of all samples are perturbed, then the quality of that dataset is considered to be 0.75. We use two noise models, one for each task. For NMT, we randomly set 50% of all non-special tokens in a selected sample to pad tokens, we do not discriminate between source and target here. For CLM, we randomly swap 50% of all non-special tokens with valid non-special tokens from the tokenizer vocabulary.
5.3 Data Sampling Strategy
We employ a nested subset sampling strategy that guarantees noise and sample monotonicity. $S=\{s_{1},$ $s_{2},$ $s_{3},\dots,$ $s_{n}\}$ is the base experiment dataset with $n$ unique samples. We draw from $S$ thrice using different seeds to create working datasets $S_{w1}$, $S_{w2}$ and $S_{w3}$. Each working dataset $S_{wi}$ is then sampled to produce datasets of different sizes. Each of these different sized datasets are then perturbed to various degrees to create the noised variants at that dataset size.
[…]
5.4 Experimental Setup
We make use of multiple compute sources in this work. NMT experiments upto a dataset volume of 1M sentence pairs were conducted on an A100 node with 4, 80GB GPUs. All other experiments are run on 2 Hopper nodes with 4, 141GB GPUs each. We do not use multi-node or multi-GPU training. We utilise the GPU counts to parallelize experiment runs.
Table 1: Estimated Parameters in Quality-Aware Scaling Law
TASK METHOD B $\bm{\beta}$ $\bm{\gamma}$ $\bm{E^{+}}$
NMT Least Squares 166.568727 0.262933 0.185135 0.146998
NMT Huber 139.602744 0.250067 0.173161 0.066539
CLM Least Squares 1428.225931 0.395142 0.388678 3.439888
CLM Huber 1441.505289 0.395859 0.400657 3.439047
Table 2: Estimated Parameters on Unseen Data - CLM
TASK METHOD B $\bm{\beta}$ $\bm{\gamma}$ $\bm{E^{+}}$
CLM (ours) Least Squares 1428.225931 0.395142 0.388678 3.439888
CLM (unseen) Least Squares 1589.071797 0.396787 0.332273 4.551611
CLM (ours) Huber 1441.505289 0.395859 0.400657 3.439047
CLM (unseen) Huber 1427.299279 0.390546 0.336753 4.540009
5.5 Main Results
(a) Machine translation
(b) Next token prediction
Figure 1: Test loss as data quality varies from 0.5 to 1 for a) machine translation and b) next token prediction.
Fitting the quality-aware scaling law.
Our main results are summarized in Table 1, which estimates $B$, $\beta$, $\gamma$, and $E^{+}$, and Figure 1, which shows the test loss as $Q$ varies from $0.5$ to $1.0$. Predictably, we find that test loss $L$ decreases as we increase the volume of data $D$ and increase the fraction of high quality samples $Q$. Table 2 shows us that using our pre-trained models to evaluate loss on unseen data also follows a scaling law similar to our fit on in distribution data, thus showing evidence for scaling generalization. To further illustrate trade-offs between $D$ and $Q$, Figure 2 presents iso-loss contours, showing how data quality improvements can substitute for increased dataset size at fixed model capacity.
Interestingly, the estimated exponents for data quality, $\hat{\gamma}$, are significantly less than one: $\hat{\gamma}\approx 0.173$ for NMT and $\hat{\gamma}\approx 0.401$ for CLM (with Huber estimation). This indicates that the effective dataset size decays sublinearly with quality, i.e., models are more robust to moderate corruption than predicted by simple effective sample-size theories from PAC learning or channel-capacity analysis, which typically suggest $\gamma\geq 1$. We hypothesize that this robustness arises from redundancy in natural language data, where even partially corrupted samples carry useful contextual information (e.g., syntax, alignment, or co-occurring clean tokens). The higher $\gamma$ in CLM compared to NMT suggests that autoregressive language modeling is more sensitive to token corruption, whereas NMT can leverage cross-sequence redundancy to mitigate the impact of noise. Another plausible hypothesis is that in the NMT noise model (padding half the tokens in a noised sample), the corrupted samples are not completely useless: alignment and context still leak enough information. In CLM (token swaps), the corruption is harsher for distributional modeling, since the noised tokens add entropy and can mislead local dependencies. Hence a larger $\gamma$.
(a) Machine translation contours
(b) Next token prediction contours
Figure 2: Iso loss contours for machine translation (a) and next token prediction (b).
Isolating the quality effect.
To quantify the incremental loss due to degraded quality at fixed $N$ and $D$, define
$\Delta L(Q)=L(N,D,Q)-L(N,D,1).$ (3)
From the fitted law, this difference takes the form
$\Delta L(Q)\;\approx\;\hat{B}\,D^{-\hat{\beta}}\,\big(Q^{-\hat{\gamma}}-1\big),$
which makes clear that $\Delta L(Q)$ should grow approximately linearly in $Q^{-\hat{\gamma}}-1$, where $\hat{\gamma}$ is the estimated $\gamma$. We therefore directly plot $\Delta L(Q)$ versus $Q^{-\hat{\gamma}}-1$, using the estimated $\hat{\gamma}$ from our scaling experiments.
As shown in Figure 3, the resulting plots are close to linear across the tested quality levels, validating that the multiplicative $Q^{-\gamma}$ term accurately captures the degradation due to corruption. Furthermore, the plots show homogeneous linear functions, which suggest that the additive terms in our proposed scaling law, $\frac{A}{N^{\alpha}}+E$, indeed do not vary with data quality $Q$. The stability of these plots across dataset sizes (0.1B, 1B, 10B for CLM; 0.5M, 1M, 2M pairs for NMT) (with the Huber estimation) further suggests that $\gamma$ is an intrinsic task-dependent parameter rather than an artifact of optimization or dataset sampling.
(a) NMT, $\hat{\gamma}=0.173$
(b) CLM, $\hat{\gamma}=0.400$
Figure 3: $\Delta L(Q)$ vs. $Q^{-\hat{\gamma}}-1$ for a) NMT and b) CLM, where $\hat{\gamma}$ is the estimated $\gamma$ parameter using the Huber estimation from the scaling experiments.
In practice, the exponent $\gamma$ serves as a robustness index: smaller values indicate that a model-task pair is resilient to corruption, while larger values reveal greater sensitivity to data quality. In our experiments, both tasks show $\gamma<1$, which means loss increases more slowly than linearly with decreasing $Q$. In other words, the model is robust to moderate amounts of corruption — one needs a large drop in $Q$ to see a noticeable rise in loss.
6 Summary and Conclusion
This paper revisits scaling laws for large language model pretraining by introducing an explicit and dimensionless measure of data quality, alongside traditional factors like model size and dataset volume. We propose a new quality-aware scaling law that extends the widely used Chinchilla framework, predicting pretraining loss as a joint function of model size, data volume, and data quality. By systematically injecting synthetic noise and varying coverage in neural machine translation and causal language modeling experiments, we demonstrate that higher data quality leads to significantly lower loss for a given model size and dataset. Notably, their findings show that for high-quality datasets, smaller models and less compute are needed to achieve strong results, which is highly relevant for domain-specific applications. The study provides a formal scaling law:
$L(N,D,Q)=\frac{A}{N^{\alpha}}+\frac{B}{D^{\beta}Q^{\gamma}}+E,$
where $Q$ measures data quality and $\gamma$ is empirically estimated. This extends earlier work that primarily focused on model and data size. Their experiments on both NMT and CLM tasks, using real datasets with systematically controlled levels of data corruption, confirm the predictive power of the new scaling law and offer concrete guidance for balancing data curation versus model scale. The work establishes a unified framework that quantifies data quality’s influence, enabling principled decisions for specialized LLM development in fields building smaller models over domain specific datasets, such as those that arise in business, scientific and medical applications.
Reproducibility Statement
We include all results from our experimental runs, all model and training configurations and recipes and all steps taken to produce this work in the relevant sections in main and the appendix.
Appendix A Proofs
A.1 Proof of Lemma 1
Proof of Lemma 1.
For the Gaussian location family, the log-likelihood of a single observation is $\ell(\theta)=-\tfrac{1}{2\sigma^{2}}(y-\theta)^{2}+\text{const}$, where $\theta=f(x)$. Differentiating twice with respect to $\theta$ and taking expectation over $\epsilon$ gives the Fisher information per sample:
$\mathcal{I}_{1}(\theta)=\mathbb{E}\!\left[-\frac{\partial^{2}}{\partial\theta^{2}}\ell(\theta)\right]=\frac{1}{\sigma^{2}}.$
By additivity of Fisher information across i.i.d. samples, $\mathcal{I}_{D}=D\cdot\mathcal{I}_{1}=D/\sigma^{2}$. Thus, the usable information is equivalent to $D_{\mathrm{eff}}=D\cdot(1/\sigma^{2})$ noise-free samples. We define the deficiency as
$\Delta=\ln\frac{\sigma^{2}}{\sigma_{0}^{2}},\qquad Q=e^{-\Delta}=\frac{\sigma_{0}^{2}}{\sigma^{2}}.$
where $\sigma_{0}^{2}$ is baseline noise (i.e. noise for “clean” data; wlog $\sigma_{0}:=1$). In this formulation, datasets noisier than the reference ($\sigma^{2}>\sigma_{0}^{2}$) have $Q<1$, while “clean” datasets ($\sigma^{2}=\sigma_{0}^{2}$) have $Q=1$. The effective sample size then scales as
$D_{\mathrm{eff}}=D\cdot Q^{\gamma}$
and taking $\gamma=1$ recovers the classical $D\cdot\mathrm{SNR}$ scaling.
A.2 Proof of Lemma 2
Proof of Lemma 2.
[…]
A.3 Proof of Proposition 1
Proof of Proposition 1.
By assumption (as in the Information Bottleneck principle [37]), the data-dependent loss scales as
$L_{D}\;\propto\;\big(D\cdot I(\tilde{X};Z)\big)^{-\beta},\qquad\beta>0.$
Substituting the multiplicative reduction of usable information,
$I(\tilde{X};Z)\;=\;\rho(Q)\,I(X;Z),$
we obtain
$L_{D}\;\propto\;D^{-\beta}\,\rho(Q)^{-\beta}\,I(X;Z)^{-\beta}.$
Since $I(X;Z)$ does not depend on $Q$ or $D$, we can absorb it into the constant $B$. By the regular-variation assumption, $\rho(Q)\approx c\,Q^{\gamma_{0}}$ as $Q\to 1$, so
$\rho(Q)^{-\beta}\;\approx\;c^{-\beta}\,Q^{-\beta\gamma_{0}}.$
Absorbing $c^{-\beta}$ into $B$ and setting $\gamma=\beta\gamma_{0}$, we obtain
$L_{D}\;\approx\;\frac{B}{D^{\beta}Q^{\gamma}}.$
Adding the standard model-size contribution $A/N^{\alpha}$ and irreducible error floor $E$ yields
$L(N,D,Q)\;\approx\;\frac{A}{N^{\alpha}}+\frac{B}{D^{\beta}Q^{\gamma}}+E,$
which is the claimed quality-aware scaling law after reparameterization of constants. ∎
Appendix B Extended Methods
B.1 Tokenizer Training
[…]
B.2 Model Configurations
We present the exact configurations that we use for our model definitions for each of the tasks. We use a GPT Neo model for NMT and a Llama model for CLM. Table 3 and Table 4 show the exact parameter definitions used for the two models. We use the official implementations exposed via HuggingFace transformers.
Variable Value
activation function gelu new
attention dropout 0.1
attention types [global, local] - 4
classifier dropout 0.1
embed dropout 0.1
hidden size 1024
initializer range 0.02
layer norm epsilon 1e-05
max position embeddings 260
num heads 8
num layers 8
resid dropout 0.1
vocab size 32000
window size 256
Table 3: GPT Neo config used for NMT pretraining experiments
Variable Value
attention bias false
attention dropout 0.0
head dim 64
hidden act silu
hidden size 512
initializer range 0.02
intermediate size 2048
max position embeddings 2048
mlp bias false
model type llama
num attention heads 8
num hidden layers 8
num key value heads 8
pretraining tp 1
rms norm eps 1e-05
rope scaling null
rope theta 10000.0
tie word embeddings true
vocab size 128256
Table 4: Llama config used for CLM pretraining experiments
B.3 Data Pre-processing
B.3.1 Paracrawl v8
In preparing our dataset for machine translation from English to German, we used a pretrained language detection model from FastText [24] lid.176.bin to filter the data as described in 5.1.
B.3.2 C4
We use a cleaned version of C4 by AllenAI [13] available on HuggingFace, they provide 5 splits, en, en.noclean, en.noblocklist, realnewslike, and multilingual (mC4). We use the en variant in our experiments, it contains 364,868,892 samples in the train split and 364,608 samples in the validation split.
Appendix C Empirical Results
C.1 Additional Experimental Details
Table 5 shows other hyperparameters and choices made during our training experiments. We also notably use greedy dataset packing in NMT and make use of DataCollatorWithFlattening() from HuggingFace in CLM to train to easily use Flash Attention 2 [11]. Additionally we use weight decay of 0.01 in both cases.
Table 5: Training Details
TASK D(avg) Batch Size Max LR Grad Acc Steps Warmup Ratio
NMT 3.6754244e+07 512 5e-4 8 0.2
NMT 7.3486728e+07 512 5e-4 8 0.2
NMT 1.4697256e+08 1024 5e-4 8 0.2
CLM 1.0317308e+08 48 1e-3 1 0.1
CLM 1.0296028e+09 48 1e-3 8 0.1
CLM 1.0294779e+10 48 1e-3 8 0.1
C.2 Additional Results
In reporting our performance we evaluate our trained models on test loss on heldout datasets. In NMT, during dataset prepartion we partition a test set of 50,000 samples and report evaluation loss on this heldout. For CLM, we partition 40,000 samples from the validation set of C4 en and use that for evaluation. The full results are shown in Table 6 and Table 7 for reference.
Table 6: Neural Machine Translation Results. $D$ is the number of tokens used for training , $Q$ represents dataset quality and $L$ is the test loss in nats. The horizontal lines demarcate the replicate experiments and vertical sections the different dataset sizes used.
D Q L
36779001 1.00 1.853
36779001 0.90 1.859
36779001 0.80 1.914
36779001 0.75 1.949
36779001 0.70 1.964
36779001 0.60 2.018
36779001 0.50 2.117
36734209 1.00 1.842
36734209 0.90 1.866
36734209 0.80 1.941
36734209 0.75 1.952
36734209 0.70 1.951
36734209 0.60 2.025
36734209 0.50 2.087
36749529 1.00 1.853
36749529 0.90 1.891
36749529 0.80 1.927
36749529 0.75 1.955
36749529 0.70 1.954
36749529 0.60 2.024
36749529 0.50 2.082
D Q L
73479862 1.00 1.565
73479862 0.90 1.596
73479862 0.80 1.629
73479862 0.75 1.662
73479862 0.70 1.672
73479862 0.60 1.715
73479862 0.50 1.799
73487922 1.00 1.572
73487922 0.90 1.602
73487922 0.80 1.631
73487922 0.75 1.655
73487922 0.70 1.688
73487922 0.60 1.722
73487922 0.50 1.791
73492417 1.00 1.576
73492417 0.90 1.607
73492417 0.80 1.626
73492417 0.75 1.662
73492417 0.70 1.673
73492417 0.60 1.715
73492417 0.50 1.791
D Q L
146952084 1.00 1.351
146952084 0.90 1.370
146952084 0.80 1.377
146952084 0.75 1.393
146952084 0.70 1.399
146952084 0.60 1.434
146952084 0.50 1.468
147005794 1.00 1.358
147005794 0.90 1.359
147005794 0.80 1.384
147005794 0.75 1.386
147005794 0.70 1.403
147005794 0.60 1.432
147005794 0.50 1.469
146959769 1.00 1.360
146959769 0.90 1.362
146959769 0.80 1.388
146959769 0.75 1.394
146959769 0.70 1.412
146959769 0.60 1.437
146959769 0.50 1.462
Table 7: Causal Language Modeling Results.DD is the number of tokens used for training , $Q$ represents dataset quality and $L$ is the test loss in nats. The horizontal lines demarcate the replicate experiments and vertical sections the different dataset sizes used.
D Q L
103068758 1.00 4.401
103068758 0.90 4.447
103068758 0.80 4.520
103068758 0.75 4.534
103068758 0.70 4.566
103068758 0.60 4.630
103068758 0.50 4.701
103036697 1.00 4.402
103036697 0.90 4.472
103036697 0.80 4.510
103036697 0.75 4.550
103036697 0.70 4.560
103036697 0.60 4.628
103036697 0.50 4.710
103413788 1.00 4.407
103413788 0.90 4.464
103413788 0.80 4.502
103413788 0.75 4.523
103413788 0.70 4.557
103413788 0.60 4.627
103413788 0.50 4.693
D Q L
1029888156 1.00 3.824
1029888156 0.90 3.846
1029888156 0.80 3.867
1029888156 0.75 3.876
1029888156 0.70 3.886
1029888156 0.60 3.917
1029888156 0.50 3.957
1029558108 1.00 3.825
1029558108 0.90 3.841
1029558108 0.80 3.866
1029558108 0.75 3.880
1029558108 0.70 3.886
1029558108 0.60 3.921
1029558108 0.50 3.950
1029362179 1.00 3.824
1029362179 0.90 3.837
1029362179 0.80 3.868
1029362179 0.75 3.878
1029362179 0.70 3.887
1029362179 0.60 3.920
1029362179 0.50 3.952
D Q L
10295030326 1.00 3.581
10295030326 0.90 3.592
10295030326 0.80 3.612
10295030326 0.75 3.621
10295030326 0.70 3.633
10295030326 0.60 3.649
10295030326 0.50 3.673
10292963774 1.00 3.584
10292963774 0.90 3.592
10292963774 0.80 3.607
10292963774 0.75 3.617
10292963774 0.70 3.629
10292963774 0.60 3.650
10292963774 0.50 3.668
10296342093 1.00 3.581
10296342093 0.90 3.589
10296342093 0.80 3.602
10296342093 0.75 3.614
10296342093 0.70 3.627
10296342093 0.60 3.646
10296342093 0.50 3.667
C.3 Parametric Loss Fitting with Huber Loss
Similar to [20] we minimize the following objective following a regularization introduced by [21]
$\min_{b,e,\beta,\gamma}\sum_{\text{Run }i}\text{Huber}_{\delta}\Big(\text{LSE}\big(b-\beta\log D_{i}-\gamma\log Q_{i},e\big)-\log L_{i}\Big)$
where $B$ = $\exp{(b)}$, $E^{+}$ = $\exp{(e)}$ and $\delta=10^{-3}$ using the L-BFGS-B solver available through SciPy minimize. We replicate the grid initialization of parameters as follows, $b\in[0,25,5]$, $e\in[0,2,0.5]$, $\beta\in[0.0,0.4,0.1]$ and $\gamma\in[0.0,0.4,0.1]$. We also add bounds $0\leq\beta\leq 1$ and $0\leq\gamma\leq 1$. We also report a secondary fit using SciPy inbuilt function curve fitting functions using Least Squares.